# Simultaneous Modeling of Protein Conformation and Dynamics via Autoregression

### v1

@@ 4 Experiments
Dataset. We evaluate model performance on ATLAS [39], a large-scale protein MD dataset covering $\sim$1300 proteins with diverse sizes and structures. For each protein, it contains triplicated 100 ns simulation trajectories. All models are trained on training trajectories and evaluated on test trajectories split by protein identity [17, 40, 18]. This presents a challenging task for assessing the generalization to unseen protein structures and dynamics.
Model training. All modules are trained from scratch, except for the FoldingModule (frozen OpenFold weights) and the DiffusionDecoder (initialized from ConfDiff). During training, sub-trajectories of length $L=8$ with varying strides ($1\sim 1024$ MD snapshots saved at 10 ps intervals) are sampled to enable learning across multiple timescales. For the main ConfRover model, we adopt a hybrid training strategy using a 1:1 ratio of trajectory and single-frame training objectives; For conformation interpolation task, we continue training the model with a 1:1:1 ratio of trajectory, single-frame and interpolation objectives. See Appendix D.2 for training details.
Baselines. We compare ConfRover with state-of-the-art deep learning models for each task: For trajectory simulation, we compare against MDGen, a flow-based non-autoregressive trajectory model trained on ATLAS; For time-independent generation, we evaluate against AlphaFlow and ConfDiff, flow- and diffusion-based conformation generation models finetuned on ATLAS; For conformation interpolation, no existing baseline is available for large proteins, so we focus on analyzing results of our model. As this is an emerging research area, baseline models remain limited. While recent works on trajectory simulation [21, 8, 6] and conformation interpolation [10, 18] have been proposed, they are not tailored for large proteins as in ATLAS or are not publicly available at the time of manuscript preparation. See Appendix A for more details.
@@ 4.1 Trajectory Simulation
Table 1: Pearson correlations of conformation changes between sampled and reference trajectories in multi-start. The better scores are highlighted in bold.
@@
Traj. Frame $\Delta$Frame
MDGen 0.55 0.45 0.40
ConfRover 0.77 0.63 0.53
@@
Traj. Frame $\Delta$Frame
MDGen 0.16 0.11 0.10
ConfRover 0.75 0.50 0.43
@@
ConfRover shows superior performance in recovering the magnitude of conformational changes. We report the Pearson correlation of measured conformation changes between model-generated and reference trajectories in Table 1, with additional results in Appendix E.1. Compare with MDGen, ConfRover shows a significant improvement in correlation scores, mean absolute error and structural quality (Table 8), indicating its stronger ability to recover the magnitude of conformation changes across different starting conditions in the conformational space. The greater difference observed in PCA highlights that ConfRover more accurately captures conformational changes along the feature dimensions most relevant to the structural variance observed in MD. Figure 4 visualizes ensembles of conformations in generated trajectories, with additional examples in Figure 12. ConfRover exhibit more notable conformation changes than MDGen, and reflects the major movements in the structured and loop domains observed in the MD reference.
MDGen is a non-autoregressive model trained on trajectories of length $L=250$. Adjusting its inference setup results in degraded conformation. To ensure fair comparison, we use the original inference setting ($S=40,L=250$) and downsampled the trajectories for evaluation. To confirm that this post-processing step does not introduce artifacts, we also trained MDGen models under the evaluation setups. The results are consistent with the downsampled version (see Appendix E.5).
@@
Table 2: State recovery in 100 ns simulation. Mean values over the first two PCA components are reported. Precision scores for all models are close to 1 and omitted. The better scores are highlighted in bold. MD 100ns is the oracle and excluded in the comparison.
@@
MDGen 0.56 0.30 0.44
ConfRover 0.51 0.42 0.58
Assessing long trajectory generation on 100 ns simulation. We further evaluate model’s ability to recover conformational states and principal dynamics of proteins. For each of protein, we simulate a trajectory of $L=80$ frames at stride $S=120$, approximating the 100 ns MD simulation in ATLAS. To assess state recovery, model-generated conformations are projected into a reduced PCA space and compared with the reference trajectory. Specifically, we discretize each principal component into 10 evenly sized “states” and measure the distribution similarity using Jensen-Shannon Distance (JSD). We also compute precision, recall, and F1-score on whether sample conformations fall within these known states [30, 40, 44]. To evaluate dynamic mode recovery, we perform time-lagged independent component analysis (tICA) at varying lag times on both reference and sample trajectories. We then compute Pearson correlations between the per-residue contribution to the leading components, based on the tICA coefficients. See Appendix E.2 for details. We include one of the triplicate MD trajectories—excluded from ground-truth evaluation—as an “oracle” reference, denoted as MD 100ns, representing the performance expected if the model were as accurate as an MD simulation run.
@@
Figure 5: Results from 100 ns simulation. (A) Correlations of principal dynamic modes between sample and reference trajectories, evaluated at varying lag time. (B) Examples trajectories illustrating the states explored by different methods (downsampled by 5 frames for visualization). The blue background indicates the density of the ground-truth conformation distribution from MD reference.
@@ 4.2 Time-independent Conformation Sampling
We evaluate the time-independent sampling performance of ConfRover, following the benchmark setup in Ye et al. [42]. For each protein, 250 independent conformations are generated and compared to MD reference ensembles. Key metrics are summarized in Table 3 with full results in Appendix E.3.
ConfRover matches the performance of state-of-the-art specialized models. Compared to AlphaFlow and ConfDiff, both specialized for time-independent conformation sampling, ConfRover achieves overall comparable performance and outperforms at least one of the SOTA models in five evaluation criteria. This demonstrate that ConfRover, despite being a general-purpose model capable of trajecotry generation, also performs strongly in sampling independent conformations that approximate the equilibrium distribution from MD simulation. In contrast, MDGen, which is trained solely for trajectory generation, shows suboptimal results with sequentially sampled conformations.
Table 3: Time-independent generation benchmark. AlphaFlow and ConfDiff are specialized SOTA models. The best scores are highlighted in bold and the second-best scores are underlined. *MDGen does not support time-independent sampling; its sequential sampling result are evaluated.
@@
AlphaFlow 0.53 0.85 2.64 1.55 2.29 0.62 0.41 0.69
ConfDiff 0.59 0.85 2.75 1.41 2.27 0.63 0.39 0.65
MDGen* 0.41 0.74 2.81 1.95 2.38 0.50 0.28 0.57
ConfRover-Traj 0.47 0.85 2.83 1.41 2.30 0.53 0.36 0.60
ConfRover 0.51 0.85 2.62 1.39 2.28 0.61 0.38 0.67
@@ 4.3 Conformation Interpolation
4.4 Extension to Masked Sequence Modeling
While the causal transformers used in ConfRover enables efficient autoregressive training and generation, proposed framework can be extended to other sequence modeling paradigm, such as masked sequence modeling. Similar to Li et al. [24], sequence models trained with bidirectional attention and masked sequence modeling can perform autoregressive generation by iteratively predicting frames in a specified order. At each iteration, the generated frame replaces its corresponding mask and joins the input sequence for the next iteration. We trained a variant ConfRover-mask following this approach. Specifically, we replaced causal attention with bidirectional attention in the transformer and implemented a scheduled masking strategy for training, linearly increasing the mask rate from 45% to 88%.
@@ 4.4 Extension to Masked Sequence Modeling
Traj. Frame $\Delta$Frame
MDGen 0.55 0.45 0.40
ConfRover 0.77 0.63 0.53
ConfRover-mask 0.71 0.60 0.49
PCA 2D
Traj. Frame $\Delta$Frame
MDGen 0.16 0.11 0.10
ConfRover 0.75 0.50 0.43
ConfRover-mask 0.73 0.49 0.41
@@ 5 Conclusions and Limitations
Nevertheless, several limitations still remain: (1) Trajectory simulation and interpolation are emerging tasks with limited baselines. We hope this work encourages further development and open-sourcing of models for comprehensive benchmarking; (2) Some evaluation metrics are still preliminary. While interpolation results are intuitive, they would benefit from more meaningful benchmarks reflecting functional state transitions. Evaluation is also constrained by the ATLAS dataset, which is limited to 100 ns simulations and mainly local dynamics. (3) Although ConfRover narrows the gap with classical MD, it still falls short in fully capturing the conformational space. Future gains may come from scaling training data, using more efficient architectures, and leveraging additional information from MD, such as energy information. (4) Finally, while the triangular updates in the structural modules ensure high accuracy, their computational cost limits scalability to larger proteins and longer trajectories. Despite these challenges, ConfRover demonstrates the promise of autoregressive models in molecular simulation, offering a unified, efficient, and extensible approach to modeling protein dynamics.
@@ 6 Acknowledgments
We thank Zaixiang Zheng for the technical discussions and valuable feedback during the preparation of this manuscript.
Appendix A Extended Discussions
@@ A.1 Related Work
Recent works have explored protein trajectory generation as a surrogate for MD simulations. Models such as Timewarp [21], ITO [34], and EquiJump [8] learn stochastic transport functions to sample future conformations at a lagged time (longer than MD intervals), reducing the computational costs of long-timescale simulations. However, these methods assume Markovian dynamics by relying solely on the current state for prediction, which may not be suitable for non-Markovian dynamics common in protein MD data. To capture higher-order dependencies between the frames, AlphaFolding [6] incorporates history frames via “motion nodes”, but it requires a fixed context window. MDGen [18] takes a different approach by directly modeling the joint distributions of frames in a trajectory and learning frame dependencies through “masked frame modeling”, similar to masked language modeling. However, its key-frame parameterization requires separate models for different tasks, and its non-autoregressive paradigm limits flexible generation (e.g., not compatible for generating trajectories with variable lengths). GST [25] applies autoregression for future frame prediction, enabling variable-length conditioning context and prediction horizons. While the autoregressive approach is conceptually similar to ConfRover, their work differs in several key aspects: it performs deterministic prediction rather than generative trajectory sampling; it employs a graph-based architecture with fixed structural priors from an adjacency graph, instead of full attention across all residues; it is trained and evaluated on a single protein instead of diverse proteins from ATLAS under a transferable setting.
Beyond forward trajectory simulation, generative models have also been applied to conformation interpolation, that is, generating the intermediate trajectories between two conditioned states. The non-autoregressive framework of MDGen [18] can be extended to sample transition pathways between such states; however, its key-frame parameterization requires training a separate model for this task, and it has not been tested on large proteins like those in ATLAS. Du et al. [10] proposed a simulation-free objective for transition-pathway sampling based on Doob’s $h$-transform, but their approach has only been validated on numerical models and the small protein Chignolin. Its generalizability to larger, more diverse proteins remains unassessed.
@@
Deep Learning Models for Molecular Conformations.
Another line of work focuses on direct sampling of conformations in a time-independent manner. Early efforts include perturbing the input to folding models (e.g., AlphaFold) [9, 41, 36] or perturbing the conformation using a structural diffusion model [30]. However, these models are trained solely on static PDB structures and do not explicitly model the conformational distribution. Recent works have shifted to deep generative paradigms that directly learn protein-specific conformational distributions [16, 44, 17, 40, 29, 23]. Several models in this category, including AlphaFlow, ConfDiff, and BioEmulator, fine-tune pretrained structure models on large-scale MD datasets, enhancing their ability to capture conformational distributions. A related approach trains normalizing flow models to approximate the Boltzmann distribution [33, 22], but their invertibility constraints limit scalability and transferability beyond small molecules and peptides. While these methods can generate time-independent conformations, they overlook temporal relationships and do not capture the kinetic aspects of protein dynamics. ConfRover handles both time-dependent trajectory generation and time-independent conformation learning.
@@
The challenge of modeling protein dynamics conceptually parallels tasks in image and video generation, requiring both data distribution learning and temporal modeling. Recent advances in video generation offer valuable insights into addressing these challenges. Given limited video data, extending image generative models to video has proven effective. Several works [15, 5, 14] achieve this by incorporating temporal attention layers, enabling frame-to-frame communication. Disabling temporal attention reverts them to image models, allowing flexible training across both modalities. These approaches efficiently model time correlations without explicitly tracking offsets between frames. Meanwhile, the extension of autoregressive language models to image and video domains has shown strong potential for sequential generation in different data modalities. Li et al. [24] integrates language models’ sequential modeling with diffusion models’ ability to model continuous distributions, showing that discrete tokens are not essential for autoregressive models. MarDini [27] extends the concept to video generation with efficient llama-style temporal planning and high-resolution video generation via a diffusion decoder. By applying masked “frame” modeling, MarDini allows the model to learn flexible temporal relationships and enables diverse tasks such as frame interpolation. ConfRover differs from these works from video models in several aspects: it employs SE(3) diffusion for 3D structure generation; by using an autoregressive paradigm, it explicitly decouples the diffusion generation from temporal modeling, unlike the spatiotemporal denoising process in MarDini; in addition, the causal autoregressive framework enables more flexible trajectory generation with variable lengths.
A.2 Training and Inference Cost
Inference cost varies with protein size and benchmark scale. For a moderately sized protein (e.g., 300 amino acids), simulating a short trajectory of 9 frames takes less than 1 minute, while generating an 80-frame trajectory (equivalent to 100 ns ATLAS setting in the paper) takes approximately 8.3 minutes. In comparison, a standard MD simulation under the ATLAS setup requires roughly 10 hours to reach 100 ns.
For our large-scale benchmarks, multi-start trajectory simulation (2,700+ trajectories) took 8 hours and 30 minutes, while time-independent sampling (250 conformations per protein) took 3 hours and 20 minutes.
A.3 Baseline Limitation
As an emerging research area, few models currently support learning protein dynamics in transferable settings. Existing approaches based on forward transport operators have been primarily trained and evaluated on small peptides (e.g., Timewarp [21]) or small fast-folding proteins (e.g., EquiJump [8]). Although AlphaFolding [6] was trained on ATLAS, its models are not yet publicly available. Due to these limitations, we use MDGen as the only available model for trajectory-based comparison. For conformation interpolation, neither sampling-based method [10] nor video-like method MDGen [18] have been trained and evaluated on large proteins.

### v2

@@ 4 Experiments
Dataset. We evaluate model performance on ATLAS [yannATLASProteinFlexibility2024], a large-scale protein MD dataset covering $\sim$1300 proteins with diverse sizes and structures. For each protein, it contains triplicated 100 ns simulation trajectories. All models are trained on training trajectories and evaluated on test trajectories split by protein identity [jingAlphaFoldMeetsFlow2024, jing2024generative_mdgen, wang2024proteinconfdiff]. This presents a challenging task for assessing the generalization to unseen protein structures and dynamics.
Model training. We use OpenFold (with frozen pretrained weights) as ConfRover’s FoldingModule and initialized the weights of the DiffusionDecoder from ConfDiff, while training the remaining parts of the network from scratch. During training, trajectories of length $L=8$ with varying timesteps (strides), corresponding to $1\sim 1024$ MD snapshots saved at 10 ps intervals, are sampled to enable learning across multiple timescales. For the base ConfRover, we adopt a hybrid training strategy with 1:1 ratio between trajectory and single-frame training objectives. To further enable conformation interpolation, we continue training the base model with a 1:1:1 ratio of trajectory, single-frame and interpolation objectives, denotes the model as ConfRover-interp. See Appendix D.2 for training details.
Baselines. We compare ConfRover with state-of-the-art deep learning models for each task: For trajectory simulation, we compare against MDGen [jing2024generative_mdgen], a flow-based non-autoregressive trajectory model trained on ATLAS; For time-independent generation, we evaluate against AlphaFlow [jingAlphaFoldMeetsFlow2024] and ConfDiff [wang2024proteinconfdiff], flow- and diffusion-based conformation generation models finetuned on ATLAS; For conformation interpolation, no existing baseline is available for large proteins, therefore, we focus on analyzing results of our model. See Appendix A for more details on the availability of baseline models.
@@ 4.1 Trajectory Simulation
Table 1: Pearson correlations between conformation changes in model-generated and reference trajectories under the multi-start setting. The mean and standard deviations are calculated from five independent runs. Models with higher correlations are highlighted in bold.
[…]
@@
[…]
ConfRover shows superior performance in recovering the magnitude of conformational changes. We report the Pearson correlation of measured conformation changes between model-generated and reference trajectories in Table 1, with additional results in Appendix E.1. Compared with MDGen, ConfRover shows a significant improvement in correlation scores, mean absolute error and structural quality (Table 9), indicating its stronger ability to recover the magnitude of conformation changes across different starting conditions in the conformational space. The greater difference observed in PCA highlights that ConfRover more accurately captures conformational changes along the feature dimensions most relevant to the structural variance observed in MD. Figure 4 visualizes ensembles of conformations in generated trajectories, with additional examples in Figure 11. ConfRover exhibit more notable conformation changes than MDGen, and reflects the major movements in the structured and loop domains observed in the MD reference.
MDGen is a non-autoregressive model trained on trajectories of length $L=250$. Adjusting its inference setup results in degraded conformation. To ensure fair comparison, we use the original inference setting ($S=40,L=250$) and downsampled the trajectories for evaluation. To confirm that this post-processing step does not introduce artifacts, we also trained MDGen models under the evaluation setups. The results are consistent with the downsampled version (see Appendix E.5.1).
@@
Table 2: Recovery of conformational states in the 100 ns simulation experiment. The mean and standard deviation are computed over five independent runs, and the better results are highlighted in bold. MD 100ns serves as the oracle and is excluded from the comparison.
@@
[…]
Assessing long trajectory generation on 100 ns simulation. We further evaluate model’s ability to recover conformational states and principal dynamics of proteins. For each of protein, we simulate a trajectory of $L=80$ frames at stride $S=120$, approximating the 100 ns MD simulation in ATLAS. To assess state recovery, model-generated conformations are projected into a reduced PCA space and compared with the reference trajectory. Specifically, we discretize each principal component into 10 evenly sized “states” and measure the distribution similarity using Jensen-Shannon Distance (JSD). We also compute precision, recall, and F1-score on whether sample conformations fall within these known states [luStr2StrScorebasedFramework2024, wang2024proteinconfdiff, zhengPredictingEquilibriumDistributions2024aDiG]. To evaluate dynamic mode recovery, we perform time-lagged independent component analysis (tICA) at varying lag times on both reference and sample trajectories. We then compute Pearson correlations between the per-residue contribution to the leading components, based on the tICA coefficients. See Appendix E.2 for details. We include one of the triplicate MD trajectories—excluded from ground-truth evaluation—as an “oracle” reference, denoted as MD 100ns, representing the performance expected if the model were as accurate as an MD simulation run.
@@
Figure 5: Results from 100 ns simulation. (A) Correlations of principal dynamic modes between sample and reference trajectories, evaluated at varying lag time. The mean and standard deviation are shown as line and shadowed area, computed from five individual runs for MDGen and ConfRover. (B) Examples trajectories illustrating the states explored by different methods (downsampled by 5 frames for visualization). The blue background indicates the density of the ground-truth conformation distribution from MD reference.
@@ 4.2 Time-independent Conformation Sampling
We evaluate the time-independent sampling performance of ConfRover, following the benchmark setup in ye2024proteinbench. For each protein, 250 independent conformations are sampled and compared the model generated ensembles with MD reference ensembles. We summarize the mean and standard deviations from five independent runs with key metrics are summarized in Table 3 with full results in Appendix E.3.
ConfRover matches the performance of state-of-the-art ensemble generation models. Compared with AlphaFlow and ConfDiff, ConfRover demonstrate overall comparable performance and outperforms at least one of the SOTA models in five evaluation criteria. This demonstrate that ConfRover, despite being a general-purpose model capable of trajectory generation, also performs strongly in sampling independent conformations that approximate the equilibrium distribution from MD simulation. In contrast, MDGen, which is trained solely for trajectory generation, shows suboptimal results with sequentially sampled conformations.
Table 3: Results from the time-independent generation experiment. AlphaFlow and ConfDiff are state-of-the-art models for direct conformation ensemble generation. ConfRover is the base model trained for both trajectory and time-independent generation, where as ConfRover-traj and MDGen are trained exclusively for trajectory generation. The mean and standard deviation are computed from five independent runs. The best scores are highlighted in bold, and the second-best scores are underlined.
@@
[…]
@@ 4.3 Conformation Interpolation
4.4 Conformation Quality
To ensure that ConfRover generates physically plausible conformations, we further evaluate the quality using geometric assessments from MolProbity package [williams2018molprobity] and energy profiles using a coarse-grained force field MadraX [orlando2024madrax]. We compared conformations across 38 trajectories shared between the forward simulation and interpolation experiments, including results from MDGen, ConfRover, and an oracle MD reference. All structures are relaxed using the refinement pipeline in OpenFold [ahdritzOpenFoldRetrainingAlphaFold22023] to enable energy comparison.
As shown in Table 4, conformations generated by MD simulation exhibit the hightest overall quality, as expected. As a generative model, ConfRover also produces high-quality conformations with fewer backbone (Ramachandran) and side-chain (rotamer) outliers, more accurate covalent lengths and angles, and achieving energy levels comparable to those of MD Reference; outperforming MDGen across all metrics. Furthermore, we compare conformations generated from forward simulation and interpolation tasks, where the latter includes additional constraints on terminal conformations. ConfRover-interp shows similar geometric and energetic metrics, indicating that intermediate conformations maintains physical plausible when ConfRover tries to interpolate between two conformational states.
Table 4: Quality of model generated conformations. Conformations sampled from 38 trajectories are evaluated. Geometric metrics are reported as mean and standard deviations across 38 cases, and energy values are reported as mean with 95% percentiles. The best scores are highlighted in bold.
Ramachandran
outliers % ($\downarrow$)
Rotamer
outliers % ($\downarrow$)
Clash
score ($\downarrow$)
RMS
bonds ($\downarrow$)
RMS
angles ($\downarrow$)
MolProbity
score ($\downarrow$)
MadraX
energy ($\downarrow$)
@@ 4.4 Conformation Quality
[…]
@@ 5 Conclusions and Limitations
Nevertheless, several limitations still remain: (1) Trajectory simulation and interpolation are emerging tasks with few available baseline models. We hope that this work, together with future developments in the field, will contribute to establishing more comprehensive benchmarks; (2) The dataset and evaluation metrics used in this study are limited and preliminary. The ATLAS dataset contains 100 ns simulations of single-chain proteins, which may not capture large conformational changes or the dynamics of protein complexes. Although we curated interpolation cases from ATLAS for demonstration purposes, future benchmarks reflecting realistic functional state transitions would be more meaningful. (3) Although ConfRover narrows the gap with classical MD, it still falls short in fully capturing the conformational space with high structural fidelity. Future gains may come from scaling training data, using more efficient architectures, and leveraging additional information from MD, such as energy information. (4) Finally, while the triangular updates in the structural modules ensure high accuracy, their computational cost limits scalability to larger proteins and longer trajectories. Despite these challenges, ConfRover demonstrates the promise of autoregressive models in molecular simulation, offering a unified, efficient, and extensible approach to modeling protein dynamics.
@@ 6 Acknowledgments
We would like to thank Dr. Hang Li for his invaluable support of this project. We also thank Zaixiang Zheng for insightful discussions, and Wesley Hsieh, Yi Zhou, Nima Shoghi, Yuxuan Liu, Xiaolu Shen, Jing Yuan, Yilai Li, Fei Ye, and Wei Qu for their valuable feedback.
Appendix A Additional Background
@@ A.1 Related Work
Recent works have explored generating protein trajectories as a surrogate for MD simulations. Models such as Timewarp [klein2024timewarp], ITO [schreiner2024implicit], and EquiJump [costa2024equijump] learn stochastic transport functions to sample future conformations at a lagged time (longer than MD intervals), reducing the computational costs of long-timescale simulations. However, these methods assume Markovian dynamics by relying solely on the current state for prediction, which may not be suitable for non-Markovian dynamics common in protein MD data. To capture higher-order dependencies between the frames, AlphaFolding [cheng20244d4ddiff] incorporates history frames via “motion nodes”, but it requires a fixed context window. MDGen [jing2024generative_mdgen] takes a different approach by directly modeling the joint distributions of frames in a trajectory and learning frame dependencies through “masked frame modeling”, similar to masked language modeling. However, its key-frame parameterization requires separate models for different tasks, and its non-autoregressive paradigm limits flexible generation (e.g., not compatible for generating trajectories with variable lengths). GST [li2025geometricGST] applies autoregression for future frame prediction, enabling variable-length conditioning context and prediction horizons. While the autoregressive approach is conceptually similar to ConfRover, their work differs in several key aspects: it performs deterministic prediction rather than generative trajectory sampling; it employs a graph-based architecture with fixed structural priors from an adjacency graph, instead of full attention across all residues; it is trained and evaluated on a single protein instead of diverse proteins from ATLAS under a transferable setting.
Beyond forward trajectory simulation, generative models have also been applied to conformation interpolation, that is, generating the intermediate trajectories between two conditioned states. The non-autoregressive framework of MDGen [jing2024generative_mdgen] can be extended to sample transition pathways between such states; however, its key-frame parameterization requires training a separate model for this task, and it has not been tested on large proteins like those in ATLAS. du2024doobs proposed a simulation-free objective for transition-pathway sampling based on Doob’s $h$-transform, but their approach has only been validated on numerical models and the small protein Chignolin. Its generalizability to larger, more diverse proteins remains unassessed.
@@
Deep Learning Models for Conformation Ensemble Generation.
Another line of work focuses on direct sampling of conformations in a time-independent manner. Early efforts include perturbing the input to folding models (e.g., AlphaFold) [delalamoSamplingAlternativeConformational2022MSAsampling, wayment-steelePredictingMultipleConformations2024MSA, steinSPEACH_AFSamplingProtein2022MSA] or perturbing the conformation using a structural diffusion model [luStr2StrScorebasedFramework2024]. However, these models are trained solely on static PDB structures and do not explicitly model the conformational distribution. Recent works have shifted to deep generative paradigms that directly learn protein-specific conformational distributions [jingEigenFoldGenerativeProtein2023EigenFold, zhengPredictingEquilibriumDistributions2024aDiG, jingAlphaFoldMeetsFlow2024, wang2024proteinconfdiff, lu2024structureesmdiff, lewis2024scalableBioemu]. Several models in this category, including AlphaFlow, ConfDiff, and BioEmulator, fine-tune pretrained structure models on large-scale MD datasets, enhancing their ability to capture conformational distributions. A related approach trains normalizing flow models to approximate the Boltzmann distribution [noe2019boltzmann, kohler2020equivariant], but their invertibility constraints limit scalability and transferability beyond small molecules and peptides. While these methods can generate time-independent conformations, they overlook temporal relationships and do not capture the kinetic aspects of protein dynamics.
Deep Learning Enhanced Molecular Dynamics.
Another direction integrates deep learning with molecular conformation modeling through machine learning force field (MLFF) models [kovacs2025maceoff, rhodes2025orbv3atomisticsimulationscale]. These models aim to incorporate higher-level accuracy (e.g., from ab initio calculations) into classical molecular dynamics simulations, improving fidelity while maintaining scalability. However, they still rely on sequential MD sampling with small integration steps and can be more computationally expensive than conventional MD. Two-for-One [Arts2023Twoforone] does not directly learn an MLFF but instead trains a diffusion model for protein conformations and uses the resulting score function as an approximate coarse-grained force field for MD simulation. Although it still depends on sequential MD sampling, this work provides an interesting perspective that connects conformation distribution modeling and trajectory generation. In contrast, different from these approaches, ConfRover explicitly models and generates both individual conformation and trajectories, without relying on force field-based MD simulations.
@@
The challenge of modeling protein dynamics conceptually parallels tasks in image and video generation, requiring both data distribution learning and temporal modeling. Recent advances in video generation offer valuable insights in addressing these challenges. Given limited video data, extending image generative models to video has proven effective. Several works [ho_video_2022, blattmann2023stablevideo, ho2022imagen] achieve this by incorporating temporal attention layers, enabling frame-to-frame communication. Disabling temporal attention reverts them to image models, allowing flexible training across both modalities. These approaches efficiently model time correlations without explicitly tracking offsets between frames. Meanwhile, the extension of autoregressive language models to image and video domains has shown strong potential for sequential generation in different data modalities. li2024autoregressivemar integrates language models’ sequential modeling with diffusion models’ ability to model continuous distributions, showing that discrete tokens are not essential for autoregressive models. MarDini [liu2024mardini] extends the concept to video generation with efficient llama-style temporal planning and high-resolution video generation via a diffusion decoder. By applying masked “frame” modeling, MarDini allows the model to learn flexible temporal relationships and enables diverse tasks such as frame interpolation. ConfRover differs from these works from video models in several aspects: it employs SE(3) diffusion for 3D structure generation; by using an autoregressive paradigm, it explicitly decouples the diffusion generation from temporal modeling, unlike the spatiotemporal denoising process in MarDini; in addition, the causal autoregressive framework enables more flexible trajectory generation with variable lengths.
A.2 Baseline Limitation
Modeling MD trajectories using deep generative models is still an emerging research area, few models currently support learning protein dynamics in transferable settings. Existing approaches based on forward transport operators have been primarily trained and evaluated on small peptides (e.g., Timewarp [klein2024timewarp]) or small fast-folding proteins (e.g., EquiJump [costa2024equijump]). Although AlphaFolding [cheng20244d4ddiff] was also trained using ATLAS, the model weights are not publicly available at the time of this work. (We attempted an internal reproduction and included its result in Appendix E.5.2.) Due to these limitations, we use MDGen as the only available model for the main experiment. For conformation interpolation, neither sampling-based method [du2024doobs] nor video-like method MDGen [jing2024generative_mdgen] have been trained and evaluated on large proteins. Therefore, we focus on demonstrating the interpolation results of ConfRover-interp.